# Communication-Efficient Module-Wise Federated Learning for Grasp Pose Detection in Cluttered Environments

### v1

@@ 6      $\Delta w_{g}\leftarrow\sum_{c=1}^{n}k_{c}\Delta w_{c}$;
10             $\Delta w_{c}[\tilde{M}]\leftarrow\texttt{ClientUpdate}(c,\tilde{M},\tau_{2})$;
11
12      $\Delta w_{g}[\tilde{M}]\leftarrow\sum_{c=1}^{n}k_{c}\Delta w_{c}[\tilde{M}]$;
@@
14       $w^{r+1}[\tilde{M}]\leftarrow w^{r+0.5}[\tilde{M}]+\Delta w_{g}[\tilde{M}]$;
15       Server distributes $w^{r+1}[\tilde{M}]$ to all clients;
16
17return $w^{R}$;
@@ IV-B Our Two-Phase Module-wise FL Algorithm
Based on our core design principle that additional effort should be focused on slower-converging modules, we introduce a novel two-phase module-wise FL algorithm. In this framework, the set of slower-converging modules, $\tilde{M}$, is pre-identified by analyzing the module-wise cosine similarity. Modules exhibiting high similarity, as discussed in Section IV-A, are selected for partial training and aggregation. Our framework, detailed in Algorithm 1, is strategically designed to resolve the fundamental trade-off between full-model updates for robust generalization and partial-model updates for communication efficiency. Each communication round combines a full-model update phase with a subsequent, efficient partial-model update phase.
Phase 1: Full-Model Training and Aggregation. As detailed in lines 1-1 of Algorithm 1, Phase 1 begins with clients training the full model ($M$) on their local data for $\tau_{1}$ epochs to compute their respective updates, $\Delta w_{c}$. These updates are then transmitted to the server, which performs a weighted aggregation to compute the global update, $\Delta w_{g}$. The server applies this to the current model ($w^{r}$) to produce an intermediate model, $w^{r+0.5}$. Finally, this intermediate model is distributed back to all clients to serve as the starting point for Phase 2.
Phase 2: Module-Selective Training and Aggregation. In the second phase, which implements partial aggregation (see Figure 1), clients use the received intermediate model $w^{r+0.5}$ as a starting point. This time, the ClientUpdate procedure is applied exclusively to the pre-identified subset of slower-converging modules, $\tilde{M}$, for $\tau_{2}$ epochs. Clients then transmit only the resulting partial updates, $\Delta w_{c}[\tilde{M}]$, back to the server. The server aggregates these to get the global partial update $\Delta w_{g}[\tilde{M}]$ and selectively applies it to the intermediate model to form the final model for the round, $w^{r+1}$ (lines 1-1). For the next round, the server then efficiently distributes only the updated partial model, $w^{r+1}[\tilde{M}]$, which clients use to reconstruct the full model locally. This approach significantly enhances communication efficiency; while Phase 1 requires a full model update, Phase 2 only transmits the parameters for the subset $\tilde{M}$, thereby reducing the overall communication.
@@ V-A Experimental Setup
Dataset and Model. Our experiments are conducted on the large-scale GraspNet-1B dataset [2], using data captured by the Kinect sensor. Our implementation builds upon the off-the-shelf codebase of EconomicGrasp [13]. We chose this framework as it achieves state-of-the-art performance among frameworks that leverage less supervision, making it an efficient baseline for our study. We adopt the standard architecture settings from the original work, with the feature dimensions of the 3D-UNet set to $[12,24,48,96,72,72,72,72]$ and a fixed learning rate of $0.001$. All other hyperparameters and data processing steps follow the original implementation unless otherwise specified. All experiments were performed on a single NVIDIA RTX 4000 GPU.
@@
Evaluation Metrics and Cost Quantification. The performance metric is Average Precision (AP), calculated across a range of friction coefficients and averaged over the top-$k$ predictions (we set $k=10$). We utilize the standard Seen, Unseen, and Novel scene splits to assess generalization. To quantify resource utilization, we define communication cost as the total volume of model parameters transmitted between the server and clients. For our proposed method, this can be expressed as:
@@
where $|w|$ is the size of the full model parameters and $|w[\tilde{M}]|$ is the size of the module-wise updates. Computation cost is the cumulative GPU processing time for a single client.
TABLE I: Quantitative comparison of our proposed method with baselines and ablation studies. We report computation costs and AP on Seen, Unseen, Novel, and Overall evaluation sets. All evaluation results are averaged over at least five runs. The best and second-best results are marked in bold and underlined, respectively.
Method Evaluation Set (AP, %) Computation Cost (sec., $\downarrow$)
@@
Baseline isolated_client $4$$\pm$ 4.50 36.57 $\pm$ 2.58 13.46 $\pm$ 1.77 31.49 $\pm$ 2.89 1000++
t_1 59.33 $\pm$ 1.43 45.57 $\pm$ 2.72 17.93 $\pm$ 1.37 40.94 $\pm$ 1.75 1642
t_2 61.70 $\pm$ 0.93 46.56 $\pm$ 1.16 18.42 $\pm$ 0.62 42.23 $\pm$ 0.69 3268
FedLAMA [25] 61.61 $\pm$ 1.23 45.35 $\pm$ 1.96 17.91 $\pm$ 1.07 41.62 $\pm$ 1.65 2996
Ablation m1 53.56 $\pm$ 4.12 34.52 $\pm$ 5.89 14.51 $\pm$ 2.62 34.19 $\pm$ 4.13 1903
m14 60.50 $\pm$ 0.64 47.30 $\pm$ 1.42 19.39 $\pm$ 1.33 42.39 $\pm$ 0.77 1754
Ours m2345 62.10 $\pm$ 0.76 48.86 $\pm$ 1.81 18.88 $\pm$ 1.30 43.28 $\pm$ 1.29 2284
m235 63.91 $\pm$ 1.17 49.82 $\pm$ 1.35 19.22 $\pm$ 1.62 44.32 $\pm$ 1.38 2588
@@ V-B Methods for Comparison
Our investigation based on the EconomicGrasp model, which we treat as a composition of five distinct modules: $m_{1}$ (3D-UNet), $m_{2}$ (GraspableNet), $m_{3}$ (ViewNet), $m_{4}$ (CGGI-Net), and $m_{5}$ (GHLI-Net). Our analysis in Section IV-A revealed that these modules exhibit heterogeneous convergence rates, with $m_{1}$ and $m_{4}$ converging faster than the others. This insight informs the design of our proposed methods and ablation studies. We compare the following approaches:
  • isolated_client (Baseline): A non-federated baseline where each client trains a model using only its own local data. There is no communication or aggregation with a server. This establishes a performance lower bound, representing the outcome without any collaboration.
  • t1 & t2 (Baselines): Standard FedAvg where all model modules are trained in each communication round. We test with the number of local epochs set to $\tau=1$ (t1) and $\tau=2$ (t2) to evaluate the impact of additional local computation.
  • FedLAMA [25] (Baseline): An existing federated learning algorithm known for its layer-wise model aggregation strategy, included for a comprehensive comparison.
  • m2345 & m235 (Ours): Our proposed module-wise FL method. These operate in a two-phase process per round: in Phase 1, all modules are trained for $\tau_{1}=1$ epoch. In Phase 2, a subset of modules $\tilde{M}$ undergoes an additional $\tau_{2}=1$ epoch of training. This strategy selectively excludes faster-converging modules. We test two variants:
    • m2345: Excludes only the backbone m1, setting $\tilde{M}=\{m_{2},m_{3},m_{4},m_{5}\}$.
    • m235: Excludes both m1 and m4, setting $\tilde{M}=\{m_{2},m_{3},m_{5}\}$.
  • m1 & m14 (Ablation): To validate our design principle, these experiments invert our strategy by applying the Phase 2 update exclusively to the fastest-converging modules, $\tilde{M}=\{m_{1}\}$ and $\tilde{M}=\{m_{1},m_{4}\}$, respectively.
@@ V-C Results and Analysis
Communication Efficiency. Figure 4 illustrates the Average Precision (AP) on the Seen evaluation set as a function of total communication cost for each method. The plot clearly indicates that our proposed methods, m235 and m2345, consistently outperform the baseline models (t1, t2, and FedLAMA) across all communication budgets. Notably, the two proposed methods exhibit distinct advantages at different stages of training. The m2345 method achieves a high level of accuracy more rapidly in the initial training phase, suggesting that its strategy is beneficial for quick initial convergence. In contrast, while starting slightly slower, m235 eventually surpasses all other methods to achieve the highest final performance. This indicates that its more focused strategy is effective for reaching peak accuracy. These findings validate that our module-wise strategy effectively utilizes a given communication budget to maximize model performance.
Quantitative Performance. Table I provides a detailed quantitative comparison of all methods. The isolated_client baseline, representing training without any collaboration, serves as a performance lower bound. Although each client model converges quickly on its small local dataset (around 1000 seconds), it achieves a low Overall AP of only 31.49%. This poor performance underscores the necessity of collaborative learning for developing robust and generalizable models.
In contrast, all federated learning methods demonstrate significant improvements. At a fixed communication cost of 7.358 GB, our proposed method, m235, achieves the highest AP across all evaluation sets, including Seen (63.91%), Unseen (49.82%), and Overall (44.32%), while our other variant, m2345, also shows strong performance by outperforming all baseline methods. Critically, m235 achieves this superior performance with a lower computational cost (2588 sec.) compared to the stronger baselines t2 (3268 sec.) and FedLAMA (2996 sec.). While the t1 baseline is computationally cheaper (1642 sec.), its performance is significantly lower. This demonstrates that our methods offer a superior trade-off between performance and computational resources.
Validation through Ablation Studies. To validate our core assertion that focusing additional training on slower-converging modules is beneficial, we conducted ablation studies (m1 and m14). In these experiments, we inverted our strategy and applied the additional training phase exclusively to the fastest-converging modules (m1 and m4, respectively). The results show that both m1 and m14 perform poorly, yielding significantly lower AP than our proposed methods and even the baselines. The performance of m1 was the lowest among all tested configurations. This outcome strongly substantiates our module selection rationale, confirming that allocating additional resources to already-converged modules is an ineffective strategy.
In summary, the experimental results robustly demonstrate that our proposed module-wise FL strategies provide a more resource-efficient path to high-performance federated GPD, striking a superior balance between model accuracy, communication efficiency, and computational expenditure.
Figure 5: The experimental setup for real-world grasp validation. The robotic platform consists of a Universal Robots UR5e arm, a Robotiq 2F-85 parallel gripper, and a wrist-mounted Intel RealSense D435 camera. The 12 numbered objects, which include a variety of common household items, were used to create cluttered test scenes.
TABLE II: Comparison of grasp success rates for different federated learning methods and object sets. The number of objects in each set is shown in parentheses.
Object Sets (# Objects) t1 m235
Attempts SR Attempts SR
1, 2, 4, 7, 8, 10 (6) 7 86% 6 100%
1, 3, 5, 7, 8, 10, 11 (7) 10 70% 9 78%
1, 3, 4, 6, 8, 9, 12 (7) 12 58% 12 58%
1, 2, 7, 8, 10, 11, 12 (7) 12 58% 14 50%
2, 3, 4, 5, 6, 7, 9, 10 (8) 15 53% 10 80%
Overall 56 62% 51 69%
@@ V-D Real-world Experiments
To further validate the practical applicability and performance of our proposed FL framework, we conducted grasping experiments on a physical robotic system. These experiments evaluated the model against inherent noise, calibration inaccuracies, and environmental variations that are not present in the simulation.
Experimental Setup. Our robotic setup consisted of a Universal Robots UR5e collaborative robot arm equipped with a Robotiq 2F-85 parallel gripper and an eye-in-hand Intel RealSense D435 depth camera for point cloud capture. A host PC running ROS2 served as the control station for communication and grasp execution. To create challenging grasping scenarios, we used a variety of common household and workshop objects, placing 6 to 8 of them in a cluttered arrangement for each experimental run. The complete setup and object set are depicted in Figure 5.
Methods and Evaluation. We compared the performance of two representative models: t1, serving as the standard FedAvg baseline, and m235, representing our proposed module-wise FL framework. The evaluation task for both models was to clear a cluttered scene of all its objects. For each run, we measured the total number of grasp attempts required to complete this task. A grasp was deemed successful if an object was securely lifted from the surface. The Success Rate (SR) for each run was then calculated as the ratio of successful grasps to the total number of attempts.
Results and Discussion. The results of our real-world experiments are summarized in Table II. The baseline t1 model achieved an overall success rate (SR) of 62% across 56 total attempts. In contrast, our proposed method, m235, demonstrated superior performance, attaining a higher overall SR of 69%. Notably, m235 was also more efficient, requiring fewer total attempts to clear all the objects in the scenes. This performance advantage indicates that the model trained with our module-wise strategy generates more reliable and effective grasp poses in a real-world environment. These experiments empirically validate that our proposed FL framework outperforms standard FedAvg not only in simulation but also in practical, physical applications, yielding a model with enhanced robustness and efficiency.

### v2

@@ 9  $\tilde{M}^{r}\leftarrow$ Select top-$\zeta$ modules with highest $S_{m}^{r}$;
12    $\Delta w_{c}[\tilde{M}^{{\color[rgb]{0,0,0}r}}]\leftarrow\texttt{ClientUpdate}(c,\tilde{M}^{{\color[rgb]{0,0,0}r}},\tau_{2})$;
13
14 $\Delta w_{g}[\tilde{M}^{{\color[rgb]{0,0,0}r}}]\leftarrow\sum_{c=1}^{n}k_{c}\Delta w_{c}[\tilde{M}^{{\color[rgb]{0,0,0}r}}]$;
@@
16 $w^{r+1}[\tilde{M}^{{\color[rgb]{0,0,0}r}}]\leftarrow w^{r+0.5}[\tilde{M}^{{\color[rgb]{0,0,0}r}}]+\Delta w_{g}[\tilde{M}^{{\color[rgb]{0,0,0}r}}]$;
17   Server distributes $w^{r+1}[\tilde{M}^{{\color[rgb]{0,0,0}r}}]$ to all clients;
18
19return $w^{R}$;
@@ IV-B Our Two-Phase Module-wise FL Algorithm
Based on our core design principle that additional effort should be focused on slower-converging modules, we introduce a novel two-phase module-wise FL algorithm. In this framework, the set of slower-converging modules, $\tilde{M}$, is adaptively identified within each communication round. This intra-round adaptation allows the algorithm to immediately focus on the modules that are the current learning bottlenecks. Our framework, detailed in Algorithm 1, is strategically designed to resolve the fundamental trade-off between full-model updates for robust generalization and partial-model updates for communication efficiency.
Phase 1: Full-Model Training and Aggregation. As detailed in lines 1-1 of Algorithm 1, each round begins with clients training the full model ($M$) on their local data for $\tau_{1}$ epochs. The resulting updates, $\Delta w_{c}$, are transmitted to the server. The server then performs a weighted aggregation to compute the global update, $\Delta w_{g}$, and applies it to the current model ($w^{r}$) to produce an intermediate model, $w^{r+0.5}$. Crucially, the full-model updates collected in this phase provide a real-time snapshot of the learning dynamics across all modules.
Phase 2: Adaptive Module-Selective Training and Aggregation. Following Phase 1, the server immediately uses the collected updates ($\Delta w_{c}$) to calculate the module-wise cosine similarity $S_{m}^{r}$ for all modules. It then identifies the top-$\zeta$ modules with the highest similarity as the current set of slow-converging modules, $\tilde{M}^{r}$ (lines 1-1). This newly identified set $\tilde{M}^{r}$, along with the intermediate model $w^{r+0.5}$, is then distributed to the clients.
In the second phase (lines 1-1), clients apply the ClientUpdate procedure exclusively to this just-in-time identified subset $\tilde{M}^{r}$ for $\tau_{2}$ epochs. Clients transmit only the resulting partial updates, $\Delta w_{c}[\tilde{M}^{r}]$, back to the server. The server aggregates these to get the global partial update $\Delta w_{g}[\tilde{M}^{{\color[rgb]{0,0,0}r}}]$ and selectively applies it to form the final model for the round, $w^{r+1}$. For the next round, the server efficiently distributes only the updated partial model, $w^{r+1}[\tilde{M}^{{\color[rgb]{0,0,0}r}}]$, which clients use to reconstruct the full model locally. This intra-round adaptive mechanism ensures that the additional training effort in Phase 2 is always directed at the most relevant modules based on the most current learning dynamics.
@@ V-A Experimental Setup
Dataset and Model. Our experiments are conducted on the large-scale GraspNet-1B dataset [fang2020graspnet], using data captured by the Kinect sensor. Our implementation builds upon the off-the-shelf codebase of EconomicGrasp [wu2024economic]. We chose this framework as it achieves state-of-the-art performance among frameworks that leverage less supervision, making it an efficient baseline for our study. We adopt the standard architecture settings from the original work, with the feature dimensions of the 3D-UNet set to $[12,24,48,96,72,72,72,72]$ and a fixed learning rate of $0.001$. All other hyperparameters and data processing steps follow the original implementation unless otherwise specified. All experiments were performed on a single NVIDIA RTX 4000 GPU.
@@
Evaluation Metrics and Cost Quantification. The performance metric is Average Precision (AP), calculated across a range of friction coefficients and averaged over the top-$k$ predictions (we set $k=10$). We utilize the standard Seen, Unseen, and Novel scene splits to assess generalization, and report an Overall score calculated as the average AP across these three splits.
To quantify resource utilization, we define communication cost as the total volume of model parameters transmitted between the server and clients. For our proposed method, this can be expressed as:
@@
where $|w|$ is the size of the full model parameters and |w[M~]||w[$\tilde${M}]| is the size of the module-wise updates. Computation cost is the cumulative GPU processing time for a single client for the entire training process.
TABLE I: Quantitative comparison of our proposed method with baselines and ablation studies. All evaluation results are averaged over at least five runs. The best evaluation results are marked in bold.
Method Evaluation Set (AP, %)
Communication
Cost (GB, $\downarrow$)
Computation
Cost (sec., $\downarrow$)
@@
Baseline iso 44.44 $\pm$ 4.50 36.57 $\pm$ 2.58 13.46 $\pm$ 1.77 31.49 $\pm$ 2.89 0 1000++
t1 62.09 $\pm$ 1.15 45.29 $\pm$ 0.61 17.92 $\pm$ 1.11 41.76 $\pm$ 0.55 10.48 2365
t2 61.72 $\pm$ 0.91 46.67 $\pm$ 1.13 18.44 $\pm$ 0.65 42.28 $\pm$ 0.67 7.417 3295
FedLAMA [lee2023layer] 62.10 $\pm$ 0.56 46.31 $\pm$ 1.45 18.59 $\pm$ 1.10 42.33 $\pm$ 0.90 7.765 3121
Ablation m1 63.20 $\pm$ 1.51 44.21 $\pm$ 3.74 17.65 $\pm$ 1.75 41.69 $\pm$ 2.24 10.41 2718
m14 62.71 $\pm$ 0.83 46.32 $\pm$ 0.71 19.08 $\pm$ 1.38 42.70 $\pm$ 0.92 9.683 2324
m35 63.68 $\pm$ 1.35 48.37 $\pm$ 1.90 19.37 $\pm$ 1.70 43.80 $\pm$ 1.64 7.947 2688
m45 64.04 $\pm$ 1.14 49.40 $\pm$ 1.01 19.34 $\pm$ 0.417 44.26 $\pm$ 0.83 7.714 2463
m135 61.75 $\pm$ 2.50 44.47 $\pm$ 1.26 17.27 $\pm$ 0.70 41.17 $\pm$ 1.46 10.44 2455
m235 63.91 $\pm$ 1.17 49.82 $\pm$ 1.35 19.22 $\pm$ 1.62 44.32 $\pm$ 1.38 7.358 2588
m245 61.69 $\pm$ 2.01 48.19 $\pm$ 2.65 18.08 $\pm$ 1.58 42.65 $\pm$ 1.84 7.717 2488
m2345 62.78 $\pm$ 0.41 49.38 $\pm$ 1.56 19.31 $\pm$ 1.27 43.82 $\pm$ 0.91 7.828 2435
Ours a3 63.17 $\pm$ 0.58 50.11 $\pm$ 1.61 19.17 $\pm$ 1.18 44.15 $\pm$ 0.69 7.358 2521
a2 64.42 $\pm$ 0.48 51.89 $\pm$ 1.64 20.59 $\pm$ 1.48 45.63 $\pm$ 0.87 6.206 2295
  • *The iso method represents isolated training and has zero communication cost by definition. Methods prefixed with m denote static ablations with pre-identified modules, while a denotes our adaptive approach where the number indicates the count of selected modules.
@@ V-B Methods for Comparison
Our investigation is based on the EconomicGrasp model, which we treat as a composition of five distinct modules: $m_{1}$ (3D-UNet), $m_{2}$ (GraspableNet), $m_{3}$ (ViewNet), $m_{4}$ (CGGI-Net), and $m_{5}$ (GHLI-Net). Our proposal is an adaptive two-phase algorithm that adaptively identifies and focuses on slow-converging modules. To validate its effectiveness, we compare it against several baselines and conduct extensive ablation studies. These studies include methods with pre-identified, static module sets for Phase 2. For all two-phase experiments, we set the local epochs to $\tau_{1}=1$ and $\tau_{2}=1$.
Baselines:
  • iso: This baseline represents an isolated client where each client trains a model using only its local data without any communication or aggregation. This establishes a performance lower bound, representing the outcome without collaboration.
  • t1 and t2: These represent standard FedAvg where the full model is trained in each communication round. We test with the number of local epochs set to $\tau=1$ (t1) and $\tau=2$ (t2) to evaluate the impact of additional local computation.
  • FedLAMA [lee2023layer]: An existing FL algorithm known for its layer-wise model aggregation strategy, included for a comprehensive comparison.
Pre-identified $\tilde{M}$ (Ablation): These experiments test our hypothesis by using various pre-identified sets of modules ($\tilde{M}$) for Phase 2. This serves to validate the strategy of targeting specific modules and motivates our adaptive approach. Each variant is denoted by m followed by the indices of the chosen modules (e.g., m14 uses the set $\tilde{M}=\{m_{1},m_{4}\}$). The combinations tested in our ablation study are m1, m14, m35, m45, m135, m235, m245, and m2345.
[…]
TABLE II: Comparison of grasp success rates for different FL methods in real-world experiments. The communication and computation costs for each method are identical to those reported in Table I.
Method Scene 1 Scene 2 Scene 3 Scene 4 Scene 5 Overall
t1 70% (10) 63% (11) 70% (10) 63% (11) 62% (13) 65% (55)
t2 88% (8) 63% (11) 63% (11) 70% (10) 73% (11) 71% (51)
m14 88% (8) 70% (10) 58% (12) 63% (11) 67% (12) 68% (53)
m235 88% (8) 70% (10) 63% (11) 63% (11) 73% (11) 71% (51)
m2345 77% (9) 77% (9) 63% (11) 70% (10) 62% (13) 69% (52)
a2 88% (8) 88% (8) 63% (11) 70% (10) 67% (12) 73% (49)
Figure 5: The experimental setup for real-world grasp validation. The robotic platform consists of a Universal Robots UR5e arm, a Robotiq 2F-85 parallel gripper, and a wrist-mounted Intel RealSense D435 camera. The 13 numbered objects, were used to create cluttered test scenes.
@@ V-C Results and Analysis
Our extensive experiments demonstrate that the proposed adaptive two-phase algorithm, particularly a2, achieves a superior balance of performance and resource efficiency for federated GPD. As summarized in Table I, our adaptive method a2 achieves the highest Overall AP of 45.63%, markedly outperforming all other approaches at a comparable communication budget. This trend is also clearly illustrated by the learning curve in Figure 4.
The Necessity of Federated Learning. First, the poor performance of the iso (isolated client) baseline, with an Overall AP of only 31.49%, confirms the necessity of FL. Without collaboration, models fail to generalize beyond their limited local data, underscoring the importance of a collaborative framework for developing robust GPD models.
Superiority Over Standard FL Baselines. Our adaptive approach, a2, showed improved performance over the standard FL baselines. With an Overall AP of 45.63%, it outperformed t1 (41.76%), t2 (42.28%), and FedLAMA (42.33%). This performance gain was also achieved with greater resource efficiency. The a2 method required 6.206 GB of communication and 2295s of computation, making it more efficient than all baselines: t1 (10.48 GB, 2365s), t2 (7.417 GB, 3295s), and FedLAMA (7.765 GB, 3121s). This result is notable as it suggests our method does not represent a typical accuracy-cost trade-off. Instead, it achieves a higher AP with a smaller resource budget, indicating that our adaptive strategy provides an efficient path for training high-performance models.
The Critical Role of Adaptive Identification. A key insight from our experiments is the importance of adaptively identifying the slower-converging modules. While several pre-identified module sets in the ablation studies, such as m45 and m235, performed well (44.26% and 44.32% Overall AP respectively), our adaptive a2 method outperformed them, achieving the highest Overall AP of 45.63%. This was also achieved with less communication overhead, as a2 required only 6.206 GB, compared to 7.714 GB for m45 and 7.358 GB for m235.
This performance gap points to a limitation of the static approach: the set of modules that act as a learning bottleneck is not fixed and can evolve during training. Static methods, which pre-identify a set $\tilde{M}$, rely on a fixed assumption about these learning dynamics. This can lead to inefficient resource allocation, as evidenced by the varied performance across the ablation studies. For instance, the m1 and m135 methods used more communication resources (10.41 GB and 10.44 GB) and similar levels of computation than the better-performing m45 (7.714 GB), yet achieved lower accuracy (41.69% and 41.17% vs. 44.26%). In contrast, our adaptive mechanism re-evaluates the learning dynamics in each communication round. This allows the additional training effort in Phase 2 to be directed toward the modules identified as most critical at that time, which explains the method’s improved efficiency and performance.
In summary, these results show that our adaptive module-wise FL strategy offers an efficient path for training high-performance GPD models in a federated setting. By adaptively focusing updates on specific modules, our method achieves higher accuracy with lower computational and communication overhead compared to both standard FL methods and the static ablation approaches.
@@ V-D Real-world Experiments
To further validate the practical applicability and performance of our proposed FL framework, we conduct grasping experiments on a physical robotic system.
Experimental Setup. Our robotic setup consisted of a Universal Robots UR5e collaborative robot arm equipped with a Robotiq 2F-85 parallel gripper and an eye-in-hand Intel RealSense D435 depth camera for point cloud capture. A host PC running ROS2 served as the control station for communication and grasp execution. To create challenging grasping scenarios, we used a variety of common household and workshop objects, placing 7 to 8 of them in a cluttered arrangement for each experimental run. The complete setup and object set are depicted in Figure 5. The object sets are as follows: Scene 1: {1, 2, 4, 9, 10, 11, 13}; Scene 2: {3, 5, 6, 7, 8, 11, 12}; Scene 3: {1, 2, 5, 8, 9, 11, 12}; Scene 4: {2, 3, 4, 5, 6, 7, 13}; Scene 5: {1, 2, 3, 4, 6, 8, 10, 13}.
Methods and Evaluation. We compared the performance of t1, t2, m14, m235, 2345, and a2. The evaluation task is to clear a cluttered scene of all its objects. For each run, we measured the total number of grasp attempts required to complete this task. A grasp was deemed successful if an object was securely lifted from the surface. The success rate for each run was then calculated by dividing the number of objects in the scene by the total number of grasp attempts required to clear them.
Results and Discussion. The results of our real-world experiments, summarized in Table II, are consistent with our simulation findings. The primary evaluation metric was the total number of grasp attempts required to clear each cluttered scene. The baseline t1 model resulted in a 65% success rate. Other methods, such as t2 and m235, performed well, requiring 51 attempts to complete the task (71% success rate). Our proposed a2 method proved to be the most effective, clearing all scenes with the fewest attempts (49), which corresponds to the highest overall success rate of 73%. This efficiency in grasp attempts suggests that the model trained with our adaptive strategy generates more reliable grasp poses, leading to fewer failures. These results support the practical applicability of our framework in a physical, cluttered environment.